# On some aspects of local thermodynamical formalism

### v2

@@ 2.2 Translocal entropy
$\overline{h}_{\omega}(z)=\max\{\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,\overline{B(z,e^{-\omega n})})\ ,\ 0\}.$
[…]
Clearly $\overline{h}_{\omega}(z)$ is decreasing in $\omega$. Compared to the quantity (2), $h_{\text{\tiny top}}(x)$ is obtained by swapping the limit $\lim_{\varepsilon\to 0}$ and $\inf_{K}$, and $\underline{h}_{\omega}(z)$ etc. comes from a simultaneous limit in between the two. As a result:
@@
Let $f:X\to X$ be a continuos map on a compact metric space. For all $x\in X$ and $\omega\geq 0$, we have
00 0 $=$ $\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\inf\left\{\frac{1}{n}\log S(n,\varepsilon,K):K\text{ is a closed $\delta$-neighbourhood of }x\right\}$
@@
$\leq$ $\inf\{h_{\text{\tiny top}}(f,K):K\text{ is a closed $\delta$-neighbourhood of }x\}\leq\ h_{\text{\tiny top}}(x).$
Proof. If $K$ is sufficiently small compared to $n$, then $S(n,\varepsilon,K)=1$, so the first equality in (2.3) is obvious. Since $S(n,\varepsilon,K)\geq 1$, we have $\frac{1}{n}\log S(n,\varepsilon,K)\geq 0$ for all $\varepsilon>0$, $n\geq 1$ and closed neighbourhoods $K$ of $x$. Also, $\frac{1}{n}\log S(n,\varepsilon,K)$ is decreasing in $n$ and in $\varepsilon$, but increasing in $K$. Write $S(\varepsilon,K)=\lim_{n\to\infty}S(n,\varepsilon,K)$; due to the subadditivity of $\log S(n,\varepsilon,K)$, the limit exists and is bounded by $h_{\text{\tiny top}}(f)$. Therefore, for every closed neighbourhood $K_{0}$ of $x$,
$0\leq\lim_{\operatorname{diam}K\to 0}S(\varepsilon,K)\leq S(\varepsilon,K_{0})\leq h_{\text{\tiny top}}(f).$
Taking the limit $\varepsilon\to 0$ on both sides gives
$0\leq\lim_{\varepsilon\to 0}\lim_{\operatorname{diam}K\to 0}S(\varepsilon,K)\leq\lim_{\varepsilon\to 0}S(\varepsilon,K_{0})\leq h_{\text{\tiny top}}(f).$
Finally take the limit $\operatorname{diam}K_{0}\to 0$, and the lemma follows. $\Box$
@@
then $h_{\text{\tiny top}}(g,K)=\log 3$, independent of the Lyapunov exponent which is non-constant for this maps. Again every point is a full entropy point. But this time, the Lyapunov exponent $\lambda(z)$ varies with the point, and it takes $k\sim\frac{\omega}{\lambda(z)}$ iterates for $K_{n}(z)$ to reach large scale. If the limit $\lambda(z)$ exists, then $h_{\omega}(z)=(1-\frac{\omega}{\lambda(z)})h_{\text{\tiny top}}(f)$.
@@
$g_{PM}(x)=\begin{cases}\frac{x}{1-x}&\text{ if }\ 0\leq x<\frac{1}{2},\\[2.84526pt] 2x-1&\text{ if }\ \frac{1}{2}\leq x<1.\end{cases}$
For this map, $g_{PM}(\frac{1}{n})=\frac{1}{n+1}$, so if $K_{n}(0)=[0,e^{-\omega n}]$, then it take $e^{\omega n}\gg n$ iterates to reach unit size. The above computation gives $h_{\omega}(0)=0$ for all $\omega>0$.
@@
This time, $g^{k}_{\sqrt{}}(e^{-\omega n})\approx 2^{k}e^{-\omega n/2^{k}}$, so it takes no more that $k\approx\log n$ iterates to reach large scale. Hence, $h_{\omega}(0)=h_{\text{\tiny top}}(g_{\sqrt{}})=\log 2$ for all $\omega\geq 0$.
@@ 2.3 Properties of translocal entropy
As the translocal entropy $h_{\omega}(z)=0$ if $f$ is constant on a neighbourhood of $z$, we will assume in the remainder of this paper that $f:X\to X$ is a local homeomorphism.
@@
The translocal entropy for iterates of the map $f$ satisfies $h_{\omega}(f^{R},z)=Rh_{\omega}(f,z)$ for integers $R\geq 0$ and also for $h_{\omega}(f^{-R},z)=Rh_{\omega}(f,z)$ if $f$ is invertible.
@@
$\limsup_{k\to\infty}\ -\frac{1}{k}\log d(f^{k}(u),v)=\infty$
Clearly $u$ has a super-exponential approach to every $v\in\operatorname{orb}_{f}(u)$, but the definition refers to other situations too. The following example shows this, but also that super-exponential approach is not a symmetric or transitive notion.
@@
Let $X=\{0,1\}^{\mathbb{Z}}$ and $\sigma:X\to X$ be the left shift. Let $w=0^{\infty}.0^{\infty}\in X$ the sequence that is constant 00. Let
@@
Note that the number of iterates $12+\sum_{j=1}^{n-1}j!-(j-1)!=(n-1!)+$ 1 before the $n$-th block of 00s is reach is negligible compared to the length $n!-(n-1)!=(n-1)(n-1)!$ of that block, so $v$ exponentially approaches $w$, but as $w$ is a fixed point, $w$ does not approach $v$. Finally, let
@@
Proof. If $\omega=0$, then the translocal entopy function coincides with the topological entropy, and there is nothing to prove. So assume that $\omega>0$.
@@
We can assume that $\operatorname{diam}f^{j}(\overline{B(w,e^{-\omega(n+k)})})<\varepsilon$ for $j\leq k$, so in the first $k$ iterates, no points in $\overline{B(w,e^{-\omega(n+k)})}$ are $\varepsilon$-separated. Therefore
@@
where in the last line we used $\frac{k^{\prime}+k}{n+k}<\frac{\log 2L}{R}$. Since $h_{\text{\tiny top}}(f)\leq\log L$ for Lipschitz maps and $\varepsilon$ and $R$ are arbitrarily, $\underline{h}_{\omega}(u)\leq\overline{h}_{\omega}(v)$, as claimed. $\Box$
@@ 2.4 Translocal entropy for one-dimensional maps
$\overline{h}_{\omega}(z)=\left(1-\omega/\overline{\lambda}(z)\right)h_{\text{\tiny top}}(f),$
where $\overline{\lambda}(z)$ is the upper Lyapunov exponent of $z$, and
$\underline{h}_{\omega}(z)=\left(1-\omega/\underline{\lambda}(z)\right)h_{\text{\tiny top}}(f),$
where $\underline{\lambda}(z)$ is the lower Lyapunov exponent of $z$.
@@
To explain the basic idea of the proof, we first assume that $f$ is a $C^{2}$ expanding circle map, i.e., there are $c,c^{+},A>0$ such that for all $x\in{\mathbb{S}}^{1}$ we have expansion: $c^{+}\geq|f^{\prime}(x)|\geq c>1$ and Adler’s condition: $A(f,x):=|f^{\prime\prime}(x)|/|f^{\prime}(x)|^{2}\leq A$ for all $x\in{\mathbb{S}}^{1}$ (see [1, 10]). This means that Adler’s conditions holds uniformly over all iterates. Indeed, due to a recursive formula $A(f\circ g,x)=(Af,g(x))+\frac{1}{f^{\prime}\circ g(x)}\cdot A(g,x)$, we get
@@
[…]
@@
In the inferior/superior limit, we get $\liminf_{n}\frac{k_{n}(z)}{n}=\frac{\omega}{\overline{\lambda}(z)}$ and $\limsup_{n}\frac{k_{n}(z)}{n}=\frac{\omega}{\underline{\lambda}(z)}$, respectively.
[…]
@@
$=$ $\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)\limsup_{n\to\infty}\frac{\log S(n-k_{n}(z)-N_{\eta},\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)-N_{\eta}}$
$\to_{\varepsilon\to 0}$ $\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)h_{\text{\tiny top}}(f).$
For the other inequality, note that restrictions $f^{j}|_{\overline{B(z,e^{-\omega n})}}$, $1\leq j\leq k_{n}(z)$, are homeomorphism, and therefore it can separate at most $1/\varepsilon$ points per iterate. Therefore
@@
$=$ $\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)\limsup_{n\to\infty}\frac{\log S(n-k_{n}(z),\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)}$
$\to_{\varepsilon\to 0}$ $\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)h_{\text{\tiny top}}(f).$
[…]
@@
$=$ $\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)\liminf_{n\to\infty}\frac{\log S(n-k_{n}(z)-N_{\eta},\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)-N_{\eta}}$
$\to_{\varepsilon\to 0}$ $\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)h_{\text{\tiny top}}(f).$
@@
$=$ $\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)\liminf_{n\to\infty}\frac{\log S(n-k_{n}(z),\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)}$
$\to_{\varepsilon\to 0}$ $\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)h_{\text{\tiny top}}(f).$
This finishes the proof. $\Box$
@@ 2.5 Translocal entropy for higher-dimensional maps
If $F=f\times g:X\to Y\to X\times Y$ is a Cartesian product and $z=(x,y)$, then $h_{\omega}(F,z)=h_{\omega}(f,x)+h_{\omega}(g,y)$.
@@
For affine toral automorphisms $F_{A}:{\mathbb{T}}^{d}\to{\mathbb{T}}^{d}$, $h_{\omega}(z)=\#\{i:\log|\lambda_{i}|\geq\omega\}(\log|\lambda_{i}|-\omega)$, where $\lambda_{i}$, $i=1,\dots,d$, are the eigenvalues of the $d\times d$-matrix $A$ that produces $F_{A}$.
Proof. Let $\mu_{1}>\mu_{2}>\dots>\mu_{k}$ be the set of absolute values of eigenvalues. Each $\mu_{j}$ represents (possibly several) eigenvalues, and the direct sums of their generalized eigenspaces is denoted as $E_{j}$; it is a hyperplane and we let $1\leq d_{j}\leq\dim(E_{j})$ be the size of the largest Jordan block in the Jordan representation of $F_{A}|_{E_{j}}$. If $K_{n,j}(z)$ is the intersection of $B(z,e^{-\omega n})$ and a local plane in the direction of $E_{j}$, then $F_{A}^{k}(K_{n,j}(z))$ reaches unit size after $k_{n}(z)=Cn^{1-d_{j}}e^{n\log\mu_{j}}$ iterates, for some uniform constant $C>0$. Here we assume that $\log\mu_{j}\geq\omega$, because otherwise $F_{A}^{k}(K_{n,j}(z))$ does not reach unit size for $k\leq n$ iterates. Therefore, $S(n,\varepsilon,K_{n,j}(z))\approx S(n-k_{n}(z),\varepsilon,E_{j}(z))$, where $E_{j}(z)$ is a unit size $\kappa_{j}$-dimensional ball centered at $z$ in the direction of $E_{j}$. Taking the logarithm and the limit $n\to\infty$, we obtain $h_{\omega}(z)=\lim_{n\to\infty}\frac{n-k_{n}(z)}{n}\frac{1}{n-k_{n}(z)}\log S(n-k_{n}(z),\varepsilon,E_{j}(z))=\kappa_{j}\max\{\log\mu_{j}-\omega,0\}$. Locally, $F_{A}$ acts as the Cartesian product of $F_{A}|_{E_{1}(z)}\times\dots\times F_{A}|_{E_{k}(z)}$. Therefore $h_{\omega}(z)=\sum_{\log\mu_{j}\geq\omega}(\log\mu_{j}-\omega)=\sum_{\log|\lambda_{i}|\geq\omega}(\log|\lambda_{i}|-\omega)$, as claimed. $\Box$
In fact, if $F$ is a toral automorphism (or Anosov map) that is not necessarily affine, but has bounded distortion, then the formula becomes $\overline{h}_{\omega}(z)=h_{\text{\tiny top}}(f)-\sum_{\log|\overline{\lambda}_{i}(z)|\geq\omega}(\log|\overline{\lambda}_{i}|(z)-\omega)$, where $\overline{\lambda}_{i}$ stands for the $i$-th upper Lyapunov exponent. The same result works for affine hyperbolic horseshoe maps (or non-affine but with bounded distortion). The global product structure of these maps is important to get this outcome, because the formula can be quite different too, as the following example shows.
@@
Let $\overline{\mathbb{D}}$ be the closed unit disk, and let $f:\overline{\mathbb{D}}\to\overline{\mathbb{D}}$ be given in polar coordinates by $f(r,\phi)\mapsto(r(2-r),3\phi\pmod{2\pi})$. The derivative of $r\mapsto r(2-r)$ at $r=0$ is $2$. Therefore it takes $K_{n}(z):=\overline{B(z,e^{-\omega n})}$ about $\frac{\omega n+\log\varepsilon}{\log 2}$ iterates to reach size $\varepsilon$, and at this point, the $\phi$-component will let $S(n,\varepsilon,K_{n}(z))$ grow, with exponential rate $\log 3$. It follows that $h_{\omega}(0)=\log 3(1-\frac{\omega}{\log 2})$.

### v3

@@ 2.2 Translocal entropy
$\overline{h}_{\omega}(z)=\lim_{\varepsilon\to 0}\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,\overline{B(z,e^{-\omega n})}).$
The lower translocal entropy $\underline{h}_{\omega}(z)$ at $z\in X$ is the same with $\liminf$ instead of $\limsup$. If the two are the same, then $h_{\omega}(z):=\overline{h}_{\omega}(z)=\underline{h}_{\omega}(z)$ is the translocal entropy at $z$.
Clearly $\overline{h}_{\omega}(z)$ is decreasing in $\omega$. Compared to the quantity (2), $h_{\text{\tiny top}}(x)$ is obtained by swapping the limit $\lim_{\varepsilon\to 0}$ and $\inf_{K}$.
As a result we get the following.
@@
Let $f:X\to X$ be a continuous map on a compact metric space. For all $x\in X$ and $\omega\geq 0$, we have
$0$ $=$ $\lim_{\varepsilon\to 0}\liminf_{n\to\infty}\inf\{\frac{1}{n}\log S(n,\varepsilon,K_{x}):\delta>0,~K_{x}\in{\mathcal{K}}_{\delta}(x)\}$
@@
$\leq$ $\inf\{h_{top}(f,K):\delta>0,~K_{x}\in{\mathcal{K}}_{\delta}(x)\}\leq\ h_{\text{\tiny top}}(x).$
Proof. If $K$ is sufficiently small compared to $n$, then $S(n,\varepsilon,K)=1$, so the first equality in (2.3) is obvious. Since $S(n,\varepsilon,K)\geq 1$, we have $\frac{1}{n}\log S(n,\varepsilon,K)\geq 0$ for all $\varepsilon>0$, $n\geq 1$ and closed neighbourhoods $K$ of $x$. Also, $\frac{1}{n}\log S(n,\varepsilon,K)$ is decreasing in $n$ and in $\varepsilon$, but increasing in $K$. Now we prove the second inequality in (2.3). To this aim we fix $n\in\mathbb{N}$ and $\varepsilon>0,$ then there exists $\delta_{n}<2e^{-\omega n}$ such that for any $K_{x}\in{\mathcal{K}}_{\delta_{n}}(x)$ the inclusion $K_{x}\subset B(x,e^{-\omega n})$ holds. Therefore, for any $n\in{\mathbb{N}}$ we get
$\frac{1}{n}\log S(n,\varepsilon,K_{x})\leq\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})).$
Taking infimum over $\delta_{n}>0$ and $K_{x}\in{\mathcal{K}}_{\delta_{n}}(x)$ we obtain
$\inf\left\{\frac{1}{n}\log S(n,\varepsilon,K_{x}):\delta_{n}>0\text{ and }K_{x}\in{\mathcal{K}}_{\delta_{n}}(x)\right\}\leq\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})).$
Finally, taking liminf $n\to\infty$ and limit $\varepsilon\to 0$ we get the second inequality in the statement of the lemma. The third inequality is obvious.
It remains to show the last inequality in (2.3). Fix $\delta>0$ and choose $K\in{\mathcal{K}}_{\delta}(x),$ then there exists $n_{K,\delta}\in\mathbb{N}$ such that for any for any $n\geq n_{K,\delta}$ the inequality $S(n,\varepsilon,B(x,e^{-\omega n})\leq S(n,\varepsilon,K)$ holds. Therefore, for any $n\geq n_{K,\delta}$
$\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})\leq\frac{1}{n}\log S(n,\varepsilon,K).$
Thus, taking infimum over $K\in{\mathcal{K}}_{\delta}(x),$ for large enough $n$ we get
$\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})\leq\inf\{\frac{1}{n}\log S(n,\varepsilon,K):~K\in{\mathcal{K}}_{\delta}(x)\},$
so
$\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})\leq\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):~K\in{\mathcal{K}}_{\delta}(x)\}$
and
$\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,B(x,e^{-\omega n})\leq\lim_{\delta\to 0}\inf\{\limsup_{n\to\infty}\frac{1}{n}\log S(n,\varepsilon,K):~K\in{\mathcal{K}}_{\delta}(x)\}.$
Finally, taking limit $\varepsilon\to 0$ we get the last inequality in (2.3). $\Box$
@@
then $h_{\text{\tiny top}}(g,K)=\log 3$, independent of the Lyapunov exponent which is non-constant for this map. Again every point is a full entropy point. But this time, the Lyapunov exponent $\lambda(z)$ varies with the point, and it takes $k\sim\frac{\omega}{\lambda(z)}$ iterates for $K_{n}(z)$ to reach large scale. If the limit $\lambda(z)$ exists, then $h_{\omega}(z)=(1-\frac{\omega}{\lambda(z)})h_{\text{\tiny top}}(f)$.
@@
$g_{PM}(x)=\begin{cases}\frac{x}{1-x}&\text{ if }\ 0\leq x\leq\frac{1}{2},\\[2.84526pt] 2x-1&\text{ if }\ \frac{1}{2}<x<1.\end{cases}$
For this map, $g_{PM}(\frac{1}{n})=\frac{1}{n-1}$ for $n\geq 2$, so if $K_{n}(0)=[0,e^{-\omega n}]$, then it takes $e^{\omega n}\gg n$ iterates to reach unit size. The above computation gives $h_{\omega}(0)=0$ for all $\omega>0$.
@@
This time, $g^{k}_{\sqrt{}}(e^{-\omega n})=2^{1-2^{-k}}e^{-\omega n2^{-k}}$, so it takes no more than $k\approx\log n$ iterates to reach large scale. Hence, $h_{\omega}(0)=h_{\text{\tiny top}}(g_{\sqrt{}})=\log 2$ for all $\omega\geq 0$.
@@ 2.3 Properties of translocal entropy
As the translocal entropy $h_{\omega}(z)=0$ if $f$ is constant on a neighbourhood of $z$, we make the assumption:
In the remainder of this paper, $f:X\to X$ is a local homeomorphism.
@@
The translocal entropy for iterates of the map $f$ satisfies $\underline{h}_{\omega}(f^{R},z)=R\underline{h}_{\omega}(f,z)$ for integers $R\geq 0$ and also for $\underline{h}_{\omega}(f^{-R},z)=R\underline{h}_{\omega}(f,z)$ for integers $R\geq 0$ if $f$ is invertible. The analogous statements hold for $\overline{h}_{\omega}$.
@@
$\limsup_{k\to\infty}\ -\frac{1}{k}\log d(f^{k}(u),v)=\infty.$
Clearly a periodic point $u$ has a super-exponential approach to every $v\in\operatorname{orb}_{f}(u)$, but the definition refers to other situations too. The following example shows this, but also that super-exponential approach is not a symmetric or transitive notion.
@@
Let $X=\{0,1\}^{\mathbb{Z}}$ and $\sigma:X\to X$ be the left shift. Let $w=0^{\infty}.0^{\infty}\in X$ the sequence that is constant 0. Let
@@
Note that the number of iterates $2+\sum_{j=1}^{n-1}j!-(j-1)!=(n-1!)+1$ before the $n$-th block of 0s is reach is negligible compared to the length n!(n1)!=(n1)(n1)!n!-(n-1)!=(n-1)(n-1)! of that block, so $v$ exponentially approaches $w$, but as $w$ is a fixed point, $w$ does not approach $v$. Finally, let
@@
Proof. If $\omega=0$, then the translocal entropy function coincides with the topological entropy, and there is nothing to prove. So assume that $\omega>0$.
@@
We can assume that $\operatorname{diam}f^{j}(\overline{B(u,e^{-\omega(n+k)})})<\varepsilon$ for $j\leq k$, so in the first $k$ iterates, no points in $\overline{B(u,e^{-\omega(n+k)})}$ are $\varepsilon$-separated. Therefore
@@
where in the last line we used $\frac{k^{\prime}+k}{n+k}<\frac{\log 2L}{R}$. Since $h_{\text{\tiny top}}(f)\leq\log L$ for Lipschitz maps and $\varepsilon$ and $R$ are arbitrary, $\underline{h}_{\omega}(u)\leq\overline{h}_{\omega}(v)$, as claimed. $\Box$
@@ 2.4 Translocal entropy for one-dimensional maps
$\overline{h}_{\omega}(z)=\max\{\left(1-\omega/\overline{\lambda}(z)\right)h_{\text{\tiny top}}(f)\ ,\ 0\},$
where $\overline{\lambda}(z)=\limsup_{n\to\infty}\frac{1}{n}\log|Df^{n}(z)|$ is the upper Lyapunov exponent of $z$, and
$\underline{h}_{\omega}(z)=\left(1-\omega/\underline{\lambda}(z)\right)h_{\text{\tiny top}}(f)\},$
where $\underline{\lambda}(z)=\liminf_{n\to\infty}\frac{1}{n}\log|Df^{n}(z)|$ is the lower Lyapunov exponent of $z$.
@@
Since $f$ is a $C^{2}$ expanding circle map, there are $c,c^{+},c_{2}>0$ such that for all $x\in{\mathbb{S}}^{1}$ we have expansion: $c^{+}\geq|f^{\prime}(x)|\geq c>1$ and $f^{\prime\prime}(x)\leq c_{2}$. Therefore Adler’s condition: $A(f,x):=|f^{\prime\prime}(x)|/|f^{\prime}(x)|^{2}\leq A:=c_{2}/c^{2}$ holds for all $x\in{\mathbb{S}}^{1}$ (see [1, 11]). This means that Adler’s conditions holds uniformly over all iterates. Indeed, due to a recursive formula $A(f\circ g,x)=(Af,g(x))+\frac{1}{f^{\prime}\circ g(x)}\cdot A(g,x)$, we get
@@
$\left|\frac{1}{(f^{k_{n}(z)})^{\prime}(\xi)}-\frac{1}{(f^{k_{n}(z)})^{\prime}(z)}\right|=\left|\int_{z}^{\xi}\left(\frac{1}{(f^{k_{n}(z)})^{\prime}(x)}\right)^{\prime}\,dx\right|=\int_{z}^{\xi}A(f^{k_{n}(z)},x)\,dx\leq\frac{Ac}{c-1}|\xi-z|.$
@@
In the inferior/superior limit, we get $\liminf_{n\to\infty}\frac{k_{n}(z)}{n}=\frac{\omega}{\overline{\lambda}(z)}$ and $\limsup_{n\to\infty}\frac{k_{n}(z)}{n}=\frac{\omega}{\underline{\lambda}(z)}$, respectively.
Let $N_{\eta}$ be such that $f^{N_{\eta}}(J)\supset{\mathbb{S}}^{1}$ for every interval $J$ of length $\eta$. Then for the statement on the $\limsup$, we get
@@
$=$ $\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)\limsup_{n\to\infty}\frac{\log S(n-k_{n}(z)-N_{\eta},\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)-N_{\eta}}.$
Taking the limit $\varepsilon\to 0$, we obtain $\overline{h}_{\omega}(z)\geq\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$.
For the other inequality, note that restrictions $f^{j}|_{\overline{B(z,e^{-\omega n})}}$, $1\leq j\leq k_{n}(z)$, are homeomorphisms, and therefore it can separate at most $1/\varepsilon$ points per iterate. Therefore
@@
$=$ $\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)\limsup_{n\to\infty}\frac{\log S(n-k_{n}(z),\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)}.\$
Taking the limit $\varepsilon\to 0$, we obtain $\overline{h}_{\omega}(z)\leq\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$, so $\overline{h}_{\omega}(z)=\left(1-\frac{\omega}{\overline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$.
Similarly, for the statement on the $\liminf$, we get
@@
$=$ $\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)\liminf_{n\to\infty}\frac{\log S(n-k_{n}(z)-N_{\eta},\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)-N_{\eta}}.$
Taking the limit $\varepsilon\to 0$, we obtain $\underline{h}_{\omega}(z)\geq\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$.
@@
$=$ $\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)\liminf_{n\to\infty}\frac{\log S(n-k_{n}(z),\varepsilon,{\mathbb{S}}^{1})}{n-k_{n}(z)}.$
Taking the limit $\varepsilon\to 0$, we obtain $\underline{h}_{\omega}(z)\leq\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$, so $\underline{h}_{\omega}(z)=\left(1-\frac{\omega}{\underline{\lambda}(z)}\right)h_{\text{\tiny top}}(f)$. This finishes the proof. $\Box$
@@ 2.5 Translocal entropy for higher-dimensional maps
If $F=f\times g:X\times Y\to X\times Y$ is a Cartesian product and $z=(x,y)$, then $h_{\omega}(F,z)=h_{\omega}(f,x)+h_{\omega}(g,y)$.
@@
Let $A$ be a unimodular integer $d\times d$-matrix and $F_{A}:{\mathbb{T}}^{d}\to{\mathbb{T}}^{d}$, $x\mapsto Ax\bmod 1$, the affine toral automorphism determined by $A$, then $h_{\omega}(z)=\sum_{\log|\lambda_{i}|\geq\omega}(\log|\lambda_{i}|-\omega)$, where $\lambda_{i}$, $i=1,\dots,d$, are the eigenvalues of the matrix $A$.
Proof. Let $\mu_{1}>\mu_{2}>\dots>\mu_{k}$ be the set of absolute values of eigenvalues and $\kappa_{j}$ is the number of eigenvalues (counted with multiplicity) of absolute value $\mu_{j}$. Each $\mu_{j}$ represents (possibly several) eigenvalues, and the direct sums of their generalized eigenspaces are denoted as $E_{j}$; it is a hyperplane and we let $1\leq d_{j}\leq\dim(E_{j})$ be the size of the largest Jordan block in the Jordan representation of $F_{A}|_{E_{j}}$. If $K_{n,j}(z)$ is the intersection of $B(z,e^{-\omega n})$ and a local plane in the direction of $E_{j}$, then $F_{A}^{k}(K_{n,j}(z))$ reaches unit size after $k_{n}(z)=Cn^{1-d_{j}}e^{n\log\mu_{j}}$ iterates, for some uniform constant $C>0$. Here we assume that $\log\mu_{j}\geq\omega$, because otherwise $F_{A}^{k}(K_{n,j}(z))$ does not reach unit size for $k\leq n$ iterates. Therefore, $S(n,\varepsilon,K_{n,j}(z))\approx S(n-k_{n}(z),\varepsilon,E_{j}(z))$, where $E_{j}(z)$ is a unit size $\kappa_{j}$-dimensional ball centered at $z$ in the direction of $E_{j}$. Taking the logarithm and the limit $n\to\infty$, we obtain $h_{\omega}(z)=\lim_{n\to\infty}\frac{n-k_{n}(z)}{n}\frac{1}{n-k_{n}(z)}\log S(n-k_{n}(z),\varepsilon,E_{j}(z))=\kappa_{j}\max\{\log\mu_{j}-\omega,0\}$. Locally, $F_{A}$ acts as the Cartesian product of $F_{A}|_{E_{1}(z)}\times\dots\times F_{A}|_{E_{k}(z)}$. Therefore $h_{\omega}(z)=\sum_{\log\mu_{j}\geq\omega}(\log\mu_{j}-\omega)=\sum_{\log|\lambda_{i}|\geq\omega}(\log|\lambda_{i}|-\omega)$, as claimed. $\Box$
In fact, if $F$ is a toral automorphism (or Anosov map) that is not necessarily affine, but has bounded distortion, then the formula becomes $\overline{h}_{\omega}(z)=h_{\text{\tiny top}}(f)-\sum_{\log|\overline{\lambda}_{i}(z)|\geq\omega}(\log|\overline{\lambda}_{i}(z)|-\omega)$, where $\overline{\lambda}_{i}$ stands for the $i$-th upper Lyapunov exponent. The same result works for affine hyperbolic horseshoe maps (or non-affine but with bounded distortion). The global product structure of these maps is important to get this outcome, because the formula can be quite different too, as the following example shows.
@@
Let $\overline{{\mathbb{D}}}$ be the closed unit disk, and let $f:\overline{{\mathbb{D}}}\to\overline{{\mathbb{D}}}$ be given in polar coordinates by $f(r,\phi)\mapsto(r(2-r),3\phi\pmod{2\pi})$. The derivative of $r\mapsto r(2-r)$ at $r=0$ is $2$. Therefore it takes $K_{n}(z):=\overline{B(z,e^{-\omega n})}$ about $\frac{\omega n+\log\varepsilon}{\log 2}$ iterates to reach size $\varepsilon$, and at this point, the $\phi$-component will let $S(n,\varepsilon,K_{n}(z))$ grow, with exponential rate $\log 3$. It follows that $h_{\omega}(0,0)=(1-\frac{\omega}{\log 2})\log 3$.